OpenStudy (anonymous):

There are 185 statistics students in a Midwestern college, and their mean age is 26.3 with a standard deviation of 8.0. If each student is required to find the ages of a sample of 16 students, what is the standard deviation of the sampling distribution of the sample means?

OpenStudy (kropot72):

The standard deviation of the sampling distribution of the sample means is found by dividing the standard deviation of the parent distribution by the square root of n, where n is the number of students in each sample. \[\sigma _{s}=\frac{\sigma}{\sqrt{n}}=\frac{8.0}{\sqrt{16}}=2.0\]

OpenStudy (anonymous):

hmm is that it? i did that and wasnt confidence!

OpenStudy (anonymous):

thank you!

OpenStudy (kropot72):

Good work on your part!!!